# Reward Is Enough: LLMs Are In-Context Reinforcement Learners

### v1

@@ Appendix A Prompt Examples
Response:
Step1: 10 - 4 = 6 (left: 6 9 13)  <Reward: 3.00>
Step2: 13 - 6 = 7 (left: 7 9)  <Reward: 0.00>
Step3: 9 * 7 = 63 (left: 63)  <Reward: 0.00>
*Answer**: (13 - (10 - 4)) * 9 = 63  <Reward: 3.00>
</attempt>
<attempt>
Input: 4 9 10 13.
Response:
Step1: 10 + 4 = 14 (left: 9 13 14)  <Reward: 0.00>
Step2: 14 + 9 = 23 (left: 13 23)  <Reward: 0.00>
Step3: 23 + 13 = 36 (left: 36)  <Reward: 0.00>
*Answer**: (10 + 4 + 9) + 13 = 36  <Reward: 0.00>
</attempt>
<attempt>
Input: 4 9 10 13.
Response:
Step1: 9 + 10 = 19 (left: 4 13 19)  <Reward: 3.00>
Step2: 19 - 13 = 6 (left: 4 6)  <Reward: 3.00>
Step3: 6 + 4 = 10 (left: 10)  <Reward: 0.00>
*Answer**: ((9 + 10) - 13) + 4 = 10  <Reward: 6.00>
</attempt>
$s_{\text{ICRL}}$
$s_{\text{task}}$
@@
Instruction: You are a seasoned text coherence evaluator. Read the TEXT below and rate its overall coherence on a scale from 1 to 10, where 1 means significantly less coherent than the Base Answer, 5 means equally coherent, and 10 means significantly more coherent. Be a strict and conservative evaluator-only assign high scores when the TEXT is clearly better than the Base Answer.
Base Answer:
@@
Rule of the Game of 24: Use all four numbers provided in the input, without repetition, and only basic arithmetic operations (+, –, ×, ÷) to obtain 24. Only three steps are allowed.
Given the following two remaining numbers from a previous step in the Game of 24, the current step is: {step}. Evaluate this step.
Examine the numbers shown in each “left: …” after the step and reason whether it is still possible to reach 24: • Sure → 3 • Likely → 1 • Impossible → 0
Return the score in the following format: **Answer**: <integer score>

### v3

@@ Appendix A Prompt Examples
Response: Step1: 10 - 4 = 6 (left: 6 9 13)  <Reward: 3.00> Step2: 13 - 6 = 7 (left: 7 9)  <Reward: 0.00> Step3: 9 * 7 = 63 (left: 63)  <Reward: 0.00> *Answer**: (13 - (10 - 4)) * 9 = 63  <Reward: 3.00> </attempt> <attempt>
Input:
4 9 10 13.
Response: Step1: 10 + 4 = 14 (left: 9 13 14)  <Reward: 0.00> Step2: 14 + 9 = 23 (left: 13 23)  <Reward: 0.00> Step3: 23 + 13 = 36 (left: 36)  <Reward: 0.00> *Answer**: (10 + 4 + 9) + 13 = 36  <Reward: 0.00> </attempt> <attempt>
Input:
4 9 10 13.
Response: Step1: 9 + 10 = 19 (left: 4 13 19)  <Reward: 3.00> Step2: 19 - 13 = 6 (left: 4 6)  <Reward: 3.00> Step3: 6 + 4 = 10 (left: 10)  <Reward: 0.00> *Answer**: ((9 + 10) - 13) + 4 = 10  <Reward: 6.00> </attempt> $s_{\text{ICRL}}$ $s_{\text{task}}$
@@
Instruction: You are a seasoned text coherence evaluator. Read the TEXT below and rate its overall coherence on a scale from 1 to 10, where 1 means significantly less coherent than the Base Answer, 5 means equally coherent, and 10 means significantly more coherent. Be a strict and conservative evaluator-only assign high scores when the TEXT is clearly better than the Base Answer. Base Answer:
@@
Rule of the Game of 24: Use all four numbers provided in the input, without repetition, and only basic arithmetic operations (+, –, ×, ÷) to obtain 24. Only three steps are allowed. Given the following two remaining numbers from a previous step in the Game of 24, the current step is: {step}. Evaluate this step. Examine the numbers shown in each “left: …” after the step and reason whether it is still possible to reach 24: • Sure → 3 • Likely → 1 • Impossible → 0 Return the score in the following format: **Answer**: <integer score>
@@ Appendix B Additional Experimental Results
B.1 Additional Benchmark: Math Competitions
Task Setup. AIME (aime) and HMMT (hmmt) are two of the most difficult mathematics competitions in the U.S. for high school students. They require not only strong mathematical knowledge but also advanced problem-solving skills. Training language models for reasoning has been one of the most effective methods for improving performance on these benchmarks (dsr1).
We aim to test whether reward signals derived from previous reasoning traces can improve subsequent reasoning attempts. To access these traces, we rely on open-source models. In addition to evaluating reasoning models on these benchmarks, we also include a broader set of open-source models to demonstrate the prevalence of ICRL capability across a wide spectrum.
In this benchmark, the model is given a single question and asked to produce an answer. The only additional instruction concerns formatting: the model must place its final answer within <answer> </answer> tags. The dataset includes the ground-truth answer for each question.
For the reward model, we use the model proposed by suCrossingRewardBridge2025, which provides a denser signal than simply checking with SymPy whether the model’s output matches the ground truth. Using such a model aligns with our belief that the growing ecosystem of judge and reward models, primarily designed for reinforcement fine-tuning of language models (lightman2023lets), will also play a key role in enabling ICRL as a test-time scaling method.
Evaluation. To verify correctness, we parse the model’s output and check whether its answer matches the ground truth answer provided in the datasets.
Baselines. We compare our method with Reflexion and Self-Refine, the two strongest baselines from the other benchmarks. Their implementations are consistent with those used in the prior experiments.
ICRL Setup. For attempts after the first, we insert the model’s previous answer (including reasoning or CoT tokens) into the context, along with its associated reward. We then prompt the model either to try a new method or to refine its best previous approaches. We truncate past answers to ensure that at least 32 prior attempts can fit in the context.
Results. Our method outperforms the baselines in almost all cases. For the reasoning mode of Qwen3-32B, our approach remains competitive with Self-Refine on AIME and surpasses it on HMMT, which is the harder benchmark. Notably, this is achieved using scalar rewards computed with three orders of magnitude less computation, while Reflexion and Self-Refine are allowed nearly twice the compute budget for generating long-CoT verbal feedback.
Appendix C Additional Analysis Results
C.1 Unseen Paper Abstract Generation
To isolate whether ICRL truly learns from external rewards rather than merely searching or selecting from the model’s parametric knowledge, we design a task where search within parametric knowledge alone is expected to be ineffective.
Setup. We fetch 30 arXiv papers published after GPT-4.1-mini’s training cutoff. Provided only the title of the papers, the model is instructed to generate the abstract. The goal is to uncover the unseen expert-written abstract as much as possible. We measure performance using ROUGE-recall between the model’s generated abstract and the ground-truth abstract, and we use the same score as the external reward. Because these abstracts lie outside the model’s pre-training distribution, success requires learning from external rewards rather than searching through memorized content.
Results. As shown in Fig. 16, Best-of-1024 sampling reaches only 0.44 ROUGE-recall, indicating that direct search over the model’s base distribution cannot recover the missing content. Self-Refine plateaus within a few rounds to 0.45 because it does not use external reward. Reflexion performs slightly better at 0.46 but largely mirrors Self-Refine, suggesting that its revisions are dominated by the model’s self-verbal feedback, which carries little useful information in this setup. In contrast, ICRL continues to improve over 200 iterations and achieves substantially higher ROUGE-recall at 0.59, demonstrating that it can effectively learn from the external reward signal and is not limited by the model’s pre-training knowledge.
Figure 16: ROUGE-Recall on Generating Unseen Paper Abstracts